OpenStudy (anonymous):

Estimate 8 1/3 x 2 4/5 using any method you like. Then find the actual product. Use your estimate to show that your answer is reasonable

OpenStudy (anonymous):

@thomaster @ladym @Preetha help me plz

OpenStudy (anonymous):

@angel12310

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

HELP! HELP! HELP! HELP!!!!!!!

hero (hero):

I prefer the expanded form \[\left(8 + \frac{1}{3}\right)\left(2 + \frac{4}{5}\right)\] \[=8\left(2 + \frac{4}{5}\right) + \frac{1}{3}\left(2 + \frac{4}{5}\right)\] \[=16 + \frac{32}{5} + \frac{2}{3} + \frac{4}{15}\] \[=16 + 6 + \frac{2}{5} + \frac{2}{3} + \frac{4}{5}\] \[=22 + \frac{6}{15} + \frac{10}{15} + \frac{4}{15}\] \[=22 + \frac{15}{15} + \frac{5}{15}\] \[=23 + \frac{1}{3}\]

OpenStudy (mathstudent55):

8 1/3 x 2 4/5 Estimate: 8 1/3 is close to 8. 2 4/5 is close to 3. 8 x 3 = 24 My estimate is 24 Actual product: \(8 \dfrac{1}{3} \times 2 \dfrac{4}{5} \) \(= \dfrac{25}{3} \times \dfrac{14}{5} \) \(= \dfrac{\cancel{5}\times 5}{3} \times \dfrac{14}{\cancel{5}} \) \(= \dfrac{70}{3} \) \(= 23 \dfrac{1}{3} \)